OpenStudy (nwolf):

9x-7y=-21 7y-9x=21 i came up with two different answers and do not know which one is correct, can someone help me figure this out.

OpenStudy (mathteacher1729):

Can you write the work you did? It is easier to spot errors this way. :)

OpenStudy (nwolf):

the first one I got 7/9 and -7/3 then I got 7/9 and -14/9

OpenStudy (anonymous):

y = 0 and x = -7/3. That's what i got through substitution. It seems to add up: \[9\left( -7/3 \right) -7\left( 0 \right) = -21 \rightarrow -63/3 = -21\]